OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = –3 and a5 = –768?

OpenStudy (2bornot2b):

\[(-3)\times(r)^5=-768\] Solve for r from the given equation

OpenStudy (anonymous):

r^5=256

OpenStudy (2bornot2b):

Your answer would be \[(-3)\times(r)^7\]

OpenStudy (anonymous):

3702

OpenStudy (anonymous):

?

OpenStudy (anonymous):

srry cant get answer for that one

OpenStudy (anonymous):

-12288

OpenStudy (anonymous):

got it!